# Towards a Playground to Democratize Experimentation and Benchmarking of AI Agents for Network Troubleshooting

### v1

@@ 1. Approach and Motivation
The natural quest for LLMs for network troubleshooting.  Given a network problem, network engineers need to undertake rather mechanical but cumbersome steps to diagnose network issues (Wang et al., 2024a; Zhou et al., 2020). These can be summarized as (i) identifying the right telemetry signals to collect, (ii) navigating dashboards and interpreting the collected data, (iii) taking corrective actions based on the insights derived from the telemetry data (iv) iterating on the previous speculating about root-causes and what-if hypothesis. Typically, network engineers need to switch on-the-fly across different monitoring configurations while network issues are ongoing and escalate. Common examples include probing the network to explain packet drops (Zhou et al., 2020; Gao et al., 2022) or refining the detection logic – e.g., request the collection of more fine-grained queue-length data to zoom into an ongoing incident (Gong et al., 2024), enable debug-level diagnostics, etc. This manual process is still complex, slow and error-prone, as it requires expert operators to reason across multiple dimensions. Furthermore, modern network telemetry paradigms over programmable data planes – such as sketches (Namkung et al., 2023; Langlet et al., 2023) and in-band network telemetry (INT) (P4.org, 2020) – have expanded the range of available measurement strategies. The increased expressiveness may introduce additional degrees of freedom and come at the cost of greater operational complexity. Thus, we observe that while programmable data planes and new telemetry techniques have enhanced operators’ visibility on the network, human intervention still remains a primary bottleneck in network triaging.
@@
Thanks to their ability to parse multimodal data and, more recently, engage in natural-language-driven reasoning (Wang et al., 2024c, a; Aykurt et al., 2024), LLMs – as a breakthrough category of AI models – hold a special promise for assisting network operators. As a result, our community has begun to explore how traditional approaches can evolve into more automated, LLM-assisted intent-based network monitoring and diagnosis solutions.
@@
Lack of holistic platforms and benchmarks.  Existing experimentation environments (Bonofiglio et al., 2018; SRL-Labs, 2020) are often limited in scope, lacking standardized and reproducible benchmarks. Moreover, modular, low-effort platforms that abstract the underlying heterogeneity of networks (e.g., cloud, edge, mobile, satellite, etc.), protocols, failure modes and tooling remain largely underexplored. Such gap poses a barrier to the development, evaluation and cross-comparison of AI-driven observability solutions (Angi et al., 2025; Habib et al., 2025; Li et al., 2025b; Moghadas et al., 2025). Recently, benchmarks for LLM agents have been proposed in the area of network configuration, such as NetConfEval (Wang et al., 2024c). This work assumes that LLM agents can be evaluated on static benchmarks, with one-shot and offline executions. While this assumption holds true for many instances of network configuration, network troubleshooting is an inherently more dynamic and interactive setting. It requires real-time feedback loops with the network, where AI agents must not only observe but also probe, react, and refine based on the evolving system conditions. Thus, it is essential to evaluate AI agents in environments that allow for such interactive, closed-loop operations, where they can dynamically adapt their strategies based on real-time telemetry and network state.
@@
We draw attention to the risk that, in the absence of standardized evaluation practices, research in this area might become increasingly fragmented and difficult to compare. We argue for a holistic, universally accepted platform that combines an extensible benchmark suite with a flexible, plug-and-play playground. We develop a preliminary Proof-of-Concept (PoC) and validate it with a ReAct agent for a toy-case failure scenario. We then outline challenges and future directions towards our vision.
@@ 2. Proof-of-Concept
In our toy-case example, a DeepSeek-R1-0528 agent tries to triage a network issue, whose root cause is a lossy link. We inject an artificial packet loss issue on the $\textsf{s1}\rightarrow\textsf{s3}s1 → s3$ link in the topology shown in Fig. 1 (emulator box), which comprises four BMv2 switches. We task the agent with (1) detecting and (2) localizing the anomaly. The agent is prompted with the operator’s intent along with the network context and available APIs – full prompt in the User line – but it is given no information about the anomaly type. The agent can access the tools illustrated in Fig. 2 (bottom) during its thought-action trajectory. The agent begins with active probing via get_reachability(), detects loss between h1 and h3, and proceeds to query port counters using bmv2_counter_read(). Based on the retrieved statistics, it successfully localizes the fault to s3. Fig. 2 illustrates the example reasoning trajectory, consisting of 15 action steps.

### v2

@@ 1. Approach and Motivation
The natural quest for LLMs for network troubleshooting.  Given a network problem, network engineers need to undertake mechanical yet cumbersome steps to diagnose and mitigate the issue (Wang et al., 2024a; Zhou et al., 2020). They can be summarized as (i) identifying the right telemetry signals to collect, (ii) navigating dashboards and interpreting the collected data, (iii) taking corrective actions based on the insights derived from the telemetry data (iv) iterating on the previous speculating about root-causes and what-if hypothesis. Typically, engineers need to switch on-the-fly across different monitoring configurations, as network issues are ongoing and escalating. Common examples include probing the network to explain packet drops (Zhou et al., 2020; Gao et al., 2022) or refining the detection logic – e.g., request the collection of more fine-grained queue-length data to zoom into an ongoing incident (Gong et al., 2024), enable debug-level diagnostics, etc. This manual process is still complex, slow and error-prone, as it requires expert operators to reason across multiple dimensions. Furthermore, modern network telemetry paradigms over programmable data planes – such as sketches (Namkung et al., 2023; Langlet et al., 2023) and in-band network telemetry (INT) (P4.org, 2020) – have expanded the range of available measurement strategies. The increased expressiveness may introduce additional degrees of freedom and come at the cost of greater operational complexity. Thus, we observe that while programmable data planes and new telemetry techniques have enhanced operators’ visibility on the network, human intervention still remains a primary bottleneck in network triaging.
@@
Thanks to their ability to parse multimodal data and, more recently, engage in natural-language-driven reasoning (Wang et al., 2024c, a; Aykurt et al., 2024), LLMs – as a breakthrough category of AI models – hold a special promise for assisting network operators. As a result, our community has begun to explore how traditional approaches can evolve into more automated, LLM-assisted intent-based network monitoring and diagnosis solutions (Angi et al., 2025; Habib et al., 2025; Li et al., 2025b; Moghadas et al., 2025).
@@
Lack of holistic platforms and benchmarks.  Nevertheless, existing experimentation environments (Bonofiglio et al., 2018; SRL-Labs, 2020) are often limited in scope, lacking standardized and reproducible benchmarks. Recently, benchmarks for LLM agents have been proposed in the area of network configuration, such as NetConfEval (Wang et al., 2024c). This work assumes that LLM agents can be evaluated on static benchmarks, with one-shot and offline executions. While this assumption holds true for many problem instances in network configuration, network troubleshooting is an inherently more dynamic and interactive problem space. It requires real-time feedback loops with the network, where AI agents must not only observe but also probe, react, and refine based on the evolving system conditions. Thus, it is essential to evaluate AI agents in environments that allow for such interactive, closed-loop operations, where they can dynamically adapt their strategies based on real-time telemetry and network state. Unfortunately, engineers wishing to experiment with AI agents for network troubleshooting must still shoulder the significant burden of developing custom infrastructure code to enable this agent-environment interaction, as no modular, low-effort platform – abstracting network environments (e.g., cloud, edge, mobile, satellite), protocols, failure modes, and tooling – exists to date.
@@
We draw attention to the risk that, in the absence of standardized evaluation practices, research in this area remains fragmented and difficult to compare. In response, we argue for a holistic, universally accepted platform that combines an extensible benchmark suite and a flexible, plug-and-play network playground – e.g., based on existing network emulators. We develop a preliminary Proof-of-Concept (PoC) and validate it with a ReAct agent for a toy-case failure scenario. We then outline challenges and future directions towards our vision.
@@ 2. Proof-of-Concept
In our toy-case example, a DeepSeek-R1-0528 agent tries to triage a network issue, whose root cause is a lossy link. We inject an artificial packet loss issue on the $\textsf{s1}\rightarrow\textsf{s3}s1 → s3$ link in the topology shown in the emulator box of Fig. 1, which comprises four BMv2 switches. We task the agent with (1) detecting and (2) localizing the anomaly. Fig. 2 (top) illustrates the agent’s reasoning trajectory, overall consisting of 15 steps. The agent is prompted with the operator’s intent (User line) and is given no clue about the anomaly root-cause. It can access the tools illustrated in Fig. 2 (bottom) to interface with the environment. It begins with active probing via get_reachability(), detects loss between h1 and h3, then proceeds to querying port counters using bmv2_counter_read(). Based on the retrieved statistics, it successfully localizes the fault to s3.